# SAMSGL: Series-Aligned Multi-Scale Graph Learning for Spatio-Temporal Forecasting

### v2

@@ IV-A1 WeatherBench
IV-A2 PeMSD4 and PeMSD8
The PeMSD4 and PeMSD8 datasets are derived from PeMS system [45] and sampled at 5-minute intervals by [18]. PeMSD4 consists of traffic flow records from 3,848 detectors located on 29 roads within San Francisco Bay Area, ranging from Jan.1, 2018 to Feb.28, 2018. PeMSD8 comprises traffic records from 1,979 detectors situated across 8 roads in San Bernardino, covering the period from Jul.1, 2016 to Aug.31, 2016. Our experiments on these two datasets center on forecasting traffic conditions one hour ahead using the most recent hour of historical data. Both the input and output sequences have a length of 12 data points.
IV-A3 METR-LA
The dataset comprises traffic records gathered from 207 sensors situated along the Los Angeles County highway. The data spans from Mar.1, 2012 to Jun.27, 2012, with a sampling interval of 5 minutes. Our prediction task involves forecasting traffic conditions for a one-hour horizon using a historical series spanning one hour. Both input and output sequences have a length of 12 data points.
IV-B Implement Details
The experiments are carried out on a single NVIDIA RTX 8000 48GB GPU. The graph network architecture is established using the torch package in Python, along with the torch-geometric package. The training phase of the model encompass 100 epochs, with a batch size set to 32. For weather forecasting tasks, we employ a data split ratio of 7:1:2 for the training, validation, and test sets, respectively [8]. For the traffic flow prediction experiments, the datasets are split in a ratio of 6:2:2 [3]. Normalization is performed on the training, testing, and validation sets by using the mean and variance of the training sets before inputting the time series into the model.
The dimension of node features after feature embedding is set to 64, and the spatial convolution channels are also set to 64, consistent with [30]. There are 4 Graph-FC blocks in the model. Each block consists of a single multi-graph convolution layer, which contains a series aligned graph convolution, a delayed graph convolution and a local graph convolution. In each series-aligned graph convolution, the 1-dimensional convolution layer employs a kernel size of 3. Subsequently, the multi-graph convolution is succeeded by 2 fully-connected layers, each yielding an output channel size of 256 [41]. Throughout the experiments, the Mean Absolute Error (MAE) is employed as the loss function, and the model is trained using the Adam optimizer. The learning schedule follows [8], which is a multi-step decay approach. Initially, the learning rate is set as 0.005, with a learning rate decay ratio of 0.05. The minimum learning rate is constrained to $1e^{-8}$. To evaluate the performance of our proposed SAMSGL, we use the MAE and the Root Mean Square Error (RMSE) metrics.
@@ IV-B Implement Details
Figure 2: Performance comparison on various spatio-temporal datasets. Panels (a), (e), (i), and (m) are the ground truths. (b-d), (f-h), and (j-l) show, in turn, the errors of RGSL, Corrformer and our SAMSGL model on wind speed, temperature and humidity datasets, respectively. (n-p) display the errors of RGSL, MGCRN and our SAMSGL model on traffic flow dataset, respectively.
@@
(b) SAGLN (Ours)
(c) Corrformer [40]
(d) RGSL [30]
Figure 3: Comparison of forecasting performance at a time step on wind dataset.
@@
IV-C Baseline Methods
We conduct a comprehensive comparison of our model against three categories models: RNN-based models combining GCN and RNN for capturing temporal and dynamic features, spatial and temporal attention models, as well as various graph structure learning models, as is shown in Table I. We compare our model with these RNN-based model: 1) GCGRU that incorporate the GCN and RNN to capture temporal and dynamic features. 2) DCRNN applied a diffusion model to model the traffic diffusion process. 3) TGCN combines combined GCN and the gated recurrent unit. 4) AGCRN proposed a node and data adaptive graph convolution. 5) CLCRN is designed to capture the meteorology dynamics and use a conditional local convolution. 6) SGP use a randomized recurrent neural network to embed the history of the input time series into high-dimensional state representations.In these utilize spatial and temporal attention models: 1) STGCN conducted a ST-Conv block to learn spatio-temporal dependencies; 2) MSTGCN captured daily, weekly and recently temporal dependencies with attention module; 3) ASTGCN use a spatial temporal attention to predict traffic flow. We compare with these graph structure learning models: 1) MGCRN use a Meta Graph Learner to learn the heterogeneity in space and time; 2) RGSL obtain the prior graph information and the implict graph information with Regularized Graph Generation. Moreover, We compare our model with Corrformer, a new global weather forecasting method with spatial multi-scale tree structure. Furthermore, MGCRN employs an RNN architecture, but its superior performance is primarily attributed to the learning of the meta graph structure. Therefore, we categorize it as a graph structure learning model. In models that necessitate the adjacency matrix as input, we construct the adjacency matrix according to distance and road connectivity between nodes. In meteorology datasets, we identify the 25 nodes with the shortest distances as neighboring nodes. In traffic datasets, the adjacency matrix is created based on road connectivity. Corrformer, which incorporates location information, leverages the latitude and longitude of nodes for embedding location features.
TABLE III: Forecasting performance across various horizons on the METR-LA dataset.
@@ IV-C Baseline Methods
  STGCN [44] 2.88 3.47 4.59
  DCRNN [24] 2.77 3.15 3.60
  AGCRN [13] 2.86 3.25 3.68
  MGCRN [16] 2.52 2.93 3.38
  SGP [50] 2.69 3.05 3.45
@@
TABLE IV: Forecasting performance across various horizons on the meteorological datasets.
  Datasets Wind Temperature Humidity Cloud_cover    Methods Horizons & Metrics 3 6 12 3 6 12 3 6 12 3 6 12 MAE MAE MAE MAE MAE MAE MAE MAE MAE MAE MAE MAE   CLCRN [8] 0.73 1.23 1.81 0.78 1.45 2.29 0.29 0.49 0.65 1.21 1.60 1.89   MGCRN [16] 0.81 1.43 2.18 0.61 1.11 1.48 0.29 0.47 0.65 1.25 1.69 2.07   RGSL [30] 0.75 1.24 1.83 0.50 0.78 1.12 0.27 0.44 0.59 1.19 1.52 1.83   Corrformer [9] 0.81 1.33 2.02 0.52 0.85 1.16 0.31 0.49 0.65 1.26 1.67 2.02   Ours 0.60 0.92 1.35 0.41 0.65 0.93 0.25 0.40 0.53 1.15 1.49 1.78
TABLE V: Component-wise analysis and ablation study on meteorology and traffic datasets.
@@
IV-D Experimental Results
Table II displays the evaluation results of our SAMSGL method, highlighting its superiority over all existing state-of-the-art AI-based approaches on meteorology datasets and comparable performance on traffic datasets. To ensure the reliability of our experiments, we carried out five repetitions with distinct random seeds (2021, 2022, 2023, 2024, 2025), following the procedures outlined in CLCRN [8], and present the mean and standard deviation of the experimental results. Notably, our SAMSGL achieves a remarkable improvement of up to 27.25% in the RMSE metric for the wind dataset. Moreover, our SAMSGL outperforms other methods in all metrics in meteorology datasets. On the traffic datasets, our method show comparable performance. Among the baseline models, RGSL and Corrformer exhibit robust performance in meteorology datasets, while MGCRN and RGSL surpass other baseline models in traffic forecasting tasks. RGSL and MGCRN derive benefits from learning the latent graph structure, whereas Corrformer demonstrates the efficiency of series alignment for considering the time delay of nodes. Our findings from Table II indicate that although CLCRN performs admirably in meteorology datasets, it faces challenges in predicting future dynamics in traffic datasets. This discrepancy can be attributed to the conditional graph convolution within CLCRN, which aims to ensure node smoothness and computes attention scores based on location similarity. These characteristics closely mimic meteorological propagation but differ significantly from the road connectivity inherent in traffic forecasting.
@@ IV-D Experimental Results
In Fig. 4, we present the predicted temporal dynamics by MGCRN, RGSL, and our proposed SAMSGL. These methods can capture temporal evolution trends effectively but tend to perform less effectively in dealing with sudden, extreme changes over time. To further assess the methods’ ability to predict temporal changes, we highlight specific examples in Fig. 5 to present a zoomed-out view of the predicted series. Fig. 5 provides a detailed visualization of the time delays between the ground truth and predicted results. A comparative analysis against other methods underscores our model’s superior performance in forecasting both the overall trend and the fluctuations within the time series. Our predictions exhibit a higher level of accuracy compared to alternative methods. Remarkably, specific attention should be drawn to the segments highlighted by purple boxes. Here, our method excels by promptly adjusting to changes in trend within the time series, displaying minimal delays. In contrast, other methods might demonstrate discrepancies in values or encounter delays in forecasting the evolving trend. This visual comparison further solidifies the efficacy of our approach in capturing intricate temporal dynamics and making more accurate predictions. To evaluate the forecasting capabilities of different methods across different horizons, we analyze the MAE of these methods with 3-rd, 6-th, and 12-th horizons. The forecasting error increase with the increase of time steps. As shown in Table III, our method exhibits stable forecasting performance in the Metr-LA dataset, with minimal variations in prediction errors across different time steps. Additionally, in Table IV, our method consistently outperforms the competition at various time steps on meteorological datasets.
@@
IV-E Ablation Studies
@@
  • w/o MG: The SAMSGL model is examined without the MG. Spatial interactions are established solely based on node locations and the distances between them.
@@
  • w/o SA-GFC (MST): The SA and the GFC are both omitted and substituted with the Multi-Component Spatial-Temporal Graph Convolution block (MST) [18].
@@
We choose not to solely replace GFC with other spatio-temporal modules due to the unique characteristics of SA. Spatio-temporal architectures like time convolution layers fusing temporal information, which obstacles the process of finding the initial points of series, while RNN-based methods serve the node feature at each time step as hidden states, rendering alignment unnecessary. Based on the above analysis and unique characteristics of SA, we choose to exclude both the SA and the GFC, instead of solely replacing GFC with other spatio-temporal modules.
As depicted in Fig. 7, the model with all components consistently delivers a strong performance at every forecasting time step onvarous datasets. Table V clearly illustrates a significant degradation in prediction accuracy when substituting the GFC with other spatio-temporal structures. Furthermore, both the SA and the MG contribute substantially to maintaining high accuracy. Overall, this analysis reaffirms the collective importance of these components in enabling our model to consistently deliver reliable forecasts.

### v3

@@ IV.1.1 WeatherBench
IV.1.2 PeMSD4 and PeMSD8
The PeMSD4 and PeMSD8 datasets are derived from PeMS system Chen et al. (2001) and sampled at 5-minute intervals by Guo et al. Guo et al. (2019). PeMSD4 consists of traffic flow records from 3,848 detectors located on 29 roads within San Francisco Bay Area, ranging from Jan. 1, 2018 to Feb. 28, 2018. PeMSD8 comprises traffic records from 1,979 detectors situated across 8 roads in San Bernardino, covering the period from Jul. 1, 2016 to Aug. 31, 2016. Our experiments on these two datasets center on forecasting traffic conditions one hour ahead using the most recent hour of historical data. Both the input and output sequences have a length of 12 data points.
IV.1.3 METR-LA
The dataset comprises traffic records gathered from 207 sensors situated along the Los Angeles County highway Jiang et al. (2023a). The data spans from Mar. 1, 2012 to Jun. 27, 2012, with a sampling interval of 5 minutes. Our prediction task involves forecasting traffic conditions for a one-hour horizon using a historical series spanning one hour. Both input and output sequences have a length of 12 data points.
IV.2 Implement Details
The experiments are carried out on a single NVIDIA RTX 8000 48GB GPU. The graph network architecture is established using the torch package in Python, along with the torch-geometric package. The training phase of the model encompass 100 epochs, with a batch size set to 32. For weather forecasting tasks, we employ a data split ratio of 7:1:2 for the training, validation, and test sets, respectively Lin et al. (2022). For the traffic flow prediction experiments, the datasets are split in a ratio of 6:2:2 Ji et al. (2023). Normalization is performed on the training, testing, and validation sets by using the mean and variance of the training sets before inputting the time series into the model.
The dimension of node features after feature embedding is set to be 64, and the spatial convolution channels are also set to be 64, consistent with RGSL Yu et al. (2022). There are 4 Graph-FC blocks in the model. Each block consists of a single multi-graph convolution layer, which contains a series aligned graph convolution, a delayed graph convolution and a local graph convolution. In each series-aligned graph convolution, the 1-dimensional convolution layer employs a kernel size of 3. Subsequently, the multi-graph convolution is succeeded by 2 fully-connected layers, each yielding an output channel size of 256 Oreshkin et al. (2019). Throughout the experiments, the Mean Absolute Error (MAE) is employed as the loss function, and the model is trained using the Adam optimizer. The learning schedule follows Lin et al. (2022), which is a multi-step decay approach. The learning rate is initially set as 0.005, with a learning rate decay ratio of 0.05. The minimum learning rate is constrained to $1e^{-8}$. To evaluate the performance of our proposed SAMSGL, we use the MAE and the Root Mean Square Error (RMSE) metrics.
@@ IV.2 Implement Details
Figure 2: Performance comparison on various spatio-temporal datasets. Panels (a), (e), (i), and (m) are the ground truths. (b-d), (f-h), and (j-l) show, in turn, the errors of RGSL, Corrformer and our SAMSGL model on wind speed, temperature and humidity datasets, respectively. (n-p) display the errors of RGSL, MGCRN and our SAMSGL model on traffic flow dataset, respectively. The blue boxes in panels (b)-(d), (f)-(h), and (j)-(l) highlight regions with significant errors.
@@
(b) SAMGSL (Ours)
(c) Corrformer Wu et al. (2021)
(d) RGSL Yu et al. (2022)
Figure 3: Comparison of forecasting performance at a time step on wind dataset. The red boxes zoom in on typical regions to illustrate the differences in predicted spatial patterns.
@@
IV.3 Baseline Methods
We conduct a comprehensive comparison of our model against two categories models: RNN-based models combining GCN and RNN for capturing temporal and dynamic features, and time-space alternation models that alternate between temporal and spatial operations. Additionally, we also present the utilization of graph structure learning, as depicted in Table 1. We compare our model with these RNN-based model: 1) GCGRU incorporates the GCN and RNN to capture temporal and dynamic features. 2) DCRNN applies a diffusion model to model the traffic diffusion process. 3) TGCN combines the GCN and the gated recurrent unit. 4) AGCRN proposes a node and data adaptive graph convolution. 5) CLCRN is designed to capture the meteorology dynamics and use a conditional local convolution. 6) SGP uses a randomized recurrent neural network to embed the history of the input time series into high-dimensional state representations. 7) MGCRN uses a Meta Graph Learner to learn the heterogeneity in space and time. 8) RGSL obtains the prior graph information and the implict graph information with Regularized Graph Generation. In these alternating temporal and spatial layers models: 1) STGCN conducts a ST-Conv block to learn spatio-temporal dependencies; 2) MSTGCN captures daily, weekly and recently temporal dependencies with attention module; 3) ASTGCN uses a spatial temporal attention to predict traffic flow. Moreover, We compare our model with Corrformer, a new global weather forecasting method with spatial multi-scale tree structure. In models that necessitate the adjacency matrix as input, we construct the adjacency matrix according to distance and road connectivity between nodes. In meteorology datasets, we identify the 25 nodes with the shortest distances as neighboring nodes. In traffic datasets, the adjacency matrix is created based on road connectivity. Corrformer, which incorporates location information, leverages the latitude and longitude of nodes for embedding location features. The RNN-based models are configured with 32 hidden units and 2 graph recurrent layers. For the time-space alternation methods, the spatial and convolutional channels are set to be 64, with 4 temporal and spatial layers, respectively. The Corrformer model maintains the same hyper-parameters as its original vision.
Table 3: Forecasting performance across various horizons on the METR-LA dataset.
@@ IV.3 Baseline Methods
STGCN Yu, Yin, and Zhu (2018) 2.88 3.47 4.59
DCRNN Li et al. (2018) 2.77 3.15 3.60
AGCRN Bai et al. (2020) 2.86 3.25 3.68
MGCRN Jiang et al. (2023a) 2.52 2.93 3.38
SGP Cini et al. (2023) 2.69 3.05 3.45
@@
Table 4: Forecasting performance across various horizons on the meteorological datasets.
Datasets Wind Temperature Humidity Cloud cover
Methods Horizons & Metrics 3 6 12 3 6 12 3 6 12 3 6 12
MAE MAE MAE MAE MAE MAE MAE MAE MAE MAE MAE MAE
CLCRN Lin et al. (2022) 0.73 1.23 1.81 0.78 1.45 2.29 0.29 0.49 0.65 1.21 1.60 1.89
MGCRN Jiang et al. (2023a) 0.81 1.43 2.18 0.61 1.11 1.48 0.29 0.47 0.65 1.25 1.69 2.07
RGSL Yu et al. (2022) 0.75 1.24 1.83 0.50 0.78 1.12 0.27 0.44 0.59 1.19 1.52 1.83
Corrformer Wu et al. (2023) 0.81 1.33 2.02 0.52 0.85 1.16 0.31 0.49 0.65 1.26 1.67 2.02
Ours 0.60 0.92 1.35 0.41 0.65 0.93 0.25 0.40 0.53 1.15 1.49 1.78
Table 5: Comonent-wise analysis and ablation study on meteorology and traffic datasets.
@@
IV.4 Experimental Results
Table 2 displays the evaluation results of our SAMSGL method, highlighting its superiority over all existing state-of-the-art AI-based approaches on meteorology datasets and comparable performance on traffic datasets. To ensure the reliability of our experiments, we carried out five repetitions with distinct random seeds (2021, 2022, 2023, 2024, 2025), following the procedures outlined in CLCRN Lin et al. (2022), and present the mean and standard deviation of the experimental results. Notably, our SAMSGL achieves a remarkable improvement of up to 27.25% in the RMSE metric for the wind dataset. Moreover, our SAMSGL outperforms other methods in all metrics in meteorology datasets. On the traffic datasets, our method show comparable performance. Among the baseline models, RGSL and Corrformer exhibit robust performance in meteorology datasets, while MGCRN and RGSL surpass other baseline models in traffic forecasting tasks. RGSL and MGCRN derive benefits from learning the latent graph structure, whereas Corrformer demonstrates the efficiency of series alignment for considering the time delay of nodes. Our findings from Table 2 indicate that although CLCRN performs admirably in meteorology datasets, it faces challenges in predicting future dynamics in traffic datasets. This discrepancy can be attributed to the conditional graph convolution within CLCRN, which aims to ensure node smoothness and computes attention scores based on location similarity. These characteristics closely mimic meteorological propagation but differ significantly from the road connectivity inherent in traffic forecasting.
@@ IV.4 Experimental Results
In Fig. 4, we present the predicted temporal dynamics by MGCRN, RGSL, and our proposed SAMSGL. These methods can capture temporal evolution trends effectively but tend to perform less effectively in dealing with sudden, extreme changes over time. To further assess the methods’ ability to predict temporal changes, we highlight specific examples in Fig. 5 to present a zoomed-out view of the predicted series. Fig. 5 provides a detailed visualization of the time delays between the ground truth and predicted results. A comparative analysis against other methods underscores our model’s superior performance in forecasting both the overall trend and the fluctuations within the time series. Our predictions exhibit a higher level of accuracy compared to alternative methods. Remarkably, specific attention should be drawn to the segments highlighted by purple boxes. Here, our method excels by promptly adjusting to changes in trend within the time series, displaying minimal delays. In contrast, other methods might demonstrate discrepancies in values or encounter delays in forecasting the evolving trend. This visual comparison further solidifies the efficacy of our approach in capturing intricate temporal dynamics and making more accurate predictions. To evaluate the forecasting capabilities of different methods across different horizons, we analyze the MAE of these methods with 3-rd, 6-th, and 12-th horizons. The forecasting error increase with the increase of time steps. As shown in Table 3, our method exhibits stable forecasting performance in the Metr-LA dataset, with minimal variations in prediction errors across different time steps. Notably, MGCRN and SGP outperform in the 3rd and 6th predictions, which can be attributed to their ability to capture local short-term dynamics through their recurrent structures. In contrast, our model simultaneously predicts results for all time steps. Additionally, in Table 4, our method consistently outperforms the competition at various time steps on meteorological datasets.
@@
Additionally, we compare the performance of selecting different nodes as the reference node for alignment. As shown in Table 6, utilizing the average series of all nodes as the reference node yields better performance, while randomly selecting a node lacks robustness.
Table 6: Forecasting performance of various reference node selection strategies.
Reference node Wind
MAE RMSE
node 1 0.92 1.43
node 100 0.91 1.42
node 1000 0.91 1.42
mean node 0.90 1.40
IV.5 Ablation Studies
@@
  • w/o MG-G: The SAMSGL model is examined without the global graphs. Spatial interactions are established solely based on node locations and the distances between them.
  • w/o MG-L: The SAMSGL model is examined without local graph. Only global graphs learned from node embedding are utilized to represent the spatial relations.
@@
  • w/o SA-GFC (MST): The SA and the GFC are both omitted and substituted with the Multi-Component Spatial-Temporal Graph Convolution block (MST) Guo et al. (2019).
@@
We don’t solely choose to replace GFC with other spatio-temporal modules due to the unique characteristics of SA. Spatio-temporal architectures like time convolution layers fusing temporal information, which obstacles the process of finding the initial points of series, while RNN-based methods serve the node feature at each time step as hidden states, rendering alignment unnecessary. Based on the above analysis and unique characteristics of SA, we choose to exclude both the SA and the GFC, instead of solely replacing GFC with other spatio-temporal modules.
As depicted in Fig. 7, the model with all components consistently delivers a strong performance at every forecasting time step onvarous datasets. Table 5 clearly illustrates a significant degradation in prediction accuracy when substituting the GFC with other spatio-temporal structures. Both local and global graphs play a crucial role in influencing prediction accuracy, underscoring the effectiveness of learning node relations across multiple scales. Additionally, the SA module contributes to maintaining high accuracy. All components demonstrate efficiency across diverse forecasting tasks, indicating the generalizability of these modules. Overall, this analysis reaffirms the collective importance of these components in enabling our model to consistently deliver reliable forecasts.
@@ V Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (62293502, 62293504, 62173147).
Data Availability Statement
The WeatherBench data employed in this paper are available from the website https://github.com/pangeo-data/WeatherBench. The PeMSD4 and PeMSD8 datasets are available from the website https://github.com/ant-research/RGSL/tree/main/data. The METR-LA dataset is available from the website https://github.com/deepkashiwa20/MegaCRN/tree/main/METRLA